# Raising the Bar: Investigating the Values of Large Language Models via Generative Evolving Testing

### v1

@@ 3.2 Joint Learning of IRT and AIG
As noted, CAT heavily relies on a difficulty-diverse and high-quality item pool, which is unfeasible with limited data, resulting in overestimated $a_{i}$ with overly simple items and vice versa (see Fig. 4). To fill this gap, GETA resorts to Automatic Item Generation (AIG) [28] to create difficulty-tailored items. Different from conventional AIG methods based on thoughtfully crafted templates, we directly leverage the generation power of LLMs to probe the value boundaries of examinees adaptively.
@@
Concretely, we denote $d\!=\!(b,c)$ for brevity, define $q_{\bm{\theta}}(a_{i}|\bm{y}_{i},\bm{d})$ as a neural Value Estimator to assess the examinee’s value conformity/alignment in terms of its response history, where $\bm{y}_{i,\cdot}\!=\!(y_{i,1},\dots,y_{i,t})$ and $\bm{d}\!=\!(d_{1},\dots,d_{n})$, and $q_{\bm{\psi}}(d_{j}|\bm{y}_{\cdot,j})$ as an Item Parameter Estimator to infer an item’s parameters from responses of diverse LLMs, where $\bm{y}_{\cdot,j}\!=\!(y_{1,j},\cdots,y_{m,j})$. Besides, an LLM-based Item Generator, $p_{\bm{\phi}}(x|d)$, is trained to generate new test items with specified difficulty, serving as self-calibrated and evolving item pool. Unlike previous work [139, 140], we use variational inference [52] for IRT estimation, which performs faster and more accurately [124, 125].
By considering $a,d$ as latent variables, we could unify IRT estimation and generator training as modelling a joint distribution $p(\bm{x},\bm{y})$. Then, an Evidence Lower BOund (ELBO) is derived as:
$\log p(\bm{x},\bm{y})$ $\geq\mathbb{E}_{q_{\bm{\theta}}(a_{i}|\bm{y}_{i,\cdot},\bm{d})q_{\bm{\psi}}(\bm{d}|\bm{y})}[\log p(\bm{y}_{i,\cdot}|a_{i},\bm{d})]\!+\!\mathbb{E}_{q_{\bm{\psi}}(\bm{d}|\bm{y})}[\log p_{\bm{\phi}}(\bm{x}|\bm{d})]\!-\!\text{KL}[q_{\bm{\psi}}(\bm{d}|\bm{y})||p(\bm{d})]$
$+\mathbb{E}_{q_{\bm{\psi}}(\bm{d}|\bm{y})}[-\text{KL}[q_{\bm{\theta}}(a_{i}|\bm{y}_{i,\cdot},\bm{d})||q(a_{i})]]=-\mathcal{L_{GI}}(\bm{\theta},\bm{\psi},\bm{\phi}),$ (3)
@@
where $q_{\bm{\psi}}(\bm{d}|\bm{y})=\prod_{j}q_{\bm{\psi}}(d_{j}|\bm{y}_{\cdot,j})$ and $q_{\bm{\theta}}(a_{i}|\bm{y}_{i,\cdot},\bm{d})$ both follow isotropic Gaussian distributions with $p(\bm{d})=\prod_{j}p(d_{j})\sim\mathcal{N}(0,1)$ and $q(a_{i})\sim\mathcal{N}(0,1)$ as priors, respectively. For $p(\bm{y}_{i,\cdot}|a_{i},\bm{d})=\prod_{j}p(y_{i,j}|a_{i},d_{j})$, we implement it directly with the IRT-2PL model in Eq.(1). In Eq.(3), the first term learns to fit item parameters and examinee value conformity to recover the original response, the second term trains the generator, e.g., a pre-trained LLM like LLaMA-2-7B, to generate item by input difficulty, and the last two terms regularize the posterior distributions of $a$ and $d$.
By minimizing $\mathcal{L_{GI}}(\theta,\omega,\phi)$ on $\mathcal{X,Y}$ collected offline, GETA not only optimizes neural IRT estimators, but also jointly trains a generator to automatically create new test items tailored to a specified $d$ in a scalable manner without any pre-defined template, mitigating data leakage problem.
@@ 3.3 Generative Evolving Testing
Our main goal is to dynamically exploit the value boundaries of the examinee LLMs. Nevertheless, trained on static $\mathcal{X}$ and $\mathcal{Y}$, the item generator still fails to cover a wide range of item difficulty, especially the unobserved $d$. To tackle this problem, we incorporate an iterative update schema.
In this case, $d$ outside the range of static data and corresponding $x$ are both unobserved. Thus, we also consider $x$ as another latent variable following [53, 131] and model the distribution of $y$:
$\log p(\bm{y})\geq\mathbb{E}_{q(\bm{x}|\bm{y})}[-\mathcal{L_{GI}}(\bm{\theta},\bm{\psi},\bm{\phi})]\!+\!H[q(\bm{x}|\bm{y})],$ (4)
where $H$ is the entropy. By further decomposing the ELBO in Eq.(3) into two parts: $-\mathcal{L_{G}}(\bm{\phi})\!=\!\mathbb{E}_{q_{\bm{\psi}}(\bm{d}|\bm{y})}[\log p_{\bm{\phi}}(\bm{x}|\bm{d})]$ and $-\mathcal{L_{I}}(\bm{\theta},\bm{\psi})$ for other terms, we have:
$\mathcal{L}(\bm{\theta},\bm{\psi},\bm{\phi})=\underbrace{\mathbb{E}_{\hat{p}(x,y)+\hat{p}(y)q(x|y)}}_{\text{Selective Generation}}[\underbrace{\mathcal{L}_{\mathcal{I}}(\bm{\theta},\bm{\psi})}_{\text{Variational IRT}}+\underbrace{\beta\mathcal{L}_{\mathcal{G}}(\bm{\phi})}_{\text{Item Generator}}]-\beta\underbrace{\mathbb{E}_{\hat{p}(y)}[H[q(x|y)]]}_{\text{Generator Regularization}},$ (5)
where $\hat{p}(y)$ is an assumed prior of $y$, which could be a uniform distribution over a broad range of difficulty, $\beta$ is a hyper-parameter to weight the generator’s iterative update, $\hat{p}(x,y)$ is the empirical distribution formed by static $\mathcal{X}$ and $\mathcal{Y}$ to train the IRT model and initialize the generator, and the last term regularizes the generator to improve the diversity of generated items, mitigating overfitting.
Algorithm 1 GETA Algorithm
Input: $\mathcal{E}$, $p_{\bm{\phi}}(x|d)$, $\{x_{j}^{0}\},\{d_{j}^{0}\}$, $T$, $k_{1}$, $k_{2}$, $\delta_{1}$, $\delta_{2}$ and $\mathcal{S}^{*}=\emptyset$
Output: $\{\hat{a}^{T}_{i}\}^{m}_{i=1}$ and the evolved $p_{\bm{\phi}}(x|d)$
1:for $i=1,2,...,m$ do
2:     Sample $y_{i,j}^{0}\!\sim\!e_{i}(y|x_{j}^{0}),\text{for each}\ x_{j}^{0}$
3:     Calculate $\hat{a}_{i}^{0}$ with $q_{\bm{\theta}}$, $\mathcal{S}_{i}^{0}\!=\!(\{y_{i,j}^{0}\},\{d_{j}^{0}\})$
4:for $t=1,2,...,T$ do
5:     for $i=1,2,...,m$ do
6:         Calculate $\hat{d}^{t}_{i}$ for $e_{i}$ with Eq.(6),
7:         Sample $x_{j}^{t}$ with Eq.(7) $j\!=\!1$ to $k_{1}$
8:         Sample $y_{i,j}^{t}\!\sim\!e_{i}(y|x_{j}^{t}),\text{each}\ x_{j}^{t},e_{i}$
9:         Calculate $d_{j}^{t}$ by $q_{\bm{\psi}}(d|\bm{y}_{\cdot,j}^{t})$
10:         for each $(x_{j}^{t},y_{\cdot,j}^{t},d^{t}_{j}$do
11:              if $|\hat{d}^{t}_{j}-d^{t}_{j}|<\delta_{1}$ then
12:                  $\mathcal{S}_{i}^{t-1}\leftarrow\mathcal{S}_{i}^{t-1}\cup\{(y_{i,j}^{t},d^{t}_{j})\}$
13:              else if $|\hat{d}_{j}^{t}-d^{t}_{j}|>\delta_{2}$ then
14:                  $S^{*}\leftarrow S^{*}\cup\{(x_{j}^{t},d^{t}_{j})\}$
15:         $S_{i}^{t}\leftarrow S_{i}^{t-1}$, Calculate $\hat{a}_{i}^{t}$ with $S_{i}^{t}$
16:     if $|S^{*}|\geq t*k_{2}$ then
17:         Optimize $\bm{\phi}$ on $S^{*}$
$\bm{\theta},\bm{\psi},\bm{\phi}$ are first optimized on $\mathcal{X},\mathcal{Y}$. Then during the evolving testing process, we solve the best-fitting difficulties according to the conformity estimated at $t$-th step as:
@@
and obtain $b^{*}=\hat{a}^{t}$ and $c$ should be as large as possible. Therefore, we directly set the expected difficulty to $\hat{a}^{t}$ and use a relatively larger $c$. Based on $b^{*}$, GETA adaptively generates new items instead of selecting ones from the static pool, called selective generation by sampling $y\sim\hat{p}(y)$ and then $x\sim q(x|y)$ with:
$q(x|y)\approx\int q_{\bm{\psi}}(d|y)p_{\bm{\phi}}(x|d)\mathbb{I}_{\mathcal{A}}(d)\text{d}d,$ (7)
@@
where $\mathbb{I}$ is the indicator function. The original Eq.(5) requires traversing all possible $d$. To achieve selective generation and obtain more targeted items, we restrict $d$ within a neighborhood of the expected $d^{*}$, i.e., $\mathcal{A}=[d^{*}-\epsilon,d^{*}+\epsilon]$.
Eq.(5) integrates IRT optimization, generator pretraining and adaptive updating, as well as selective generation as a unified response modelling, which explores the value vulnerabilities, boundary items and conformity limitations for each LLM. Once the estimators and generator are optimized on static data, GETA starts the evolving testing of all examinees. Given few seed items $\{x^{0}_{j},d_{j}^{0}\}$, if the LLM easily responds correctly to them, GETA generates $k_{1}$ new items with a tailored difficulty $\hat{d}$, avoiding data leakage. The generated items needs to be responded by all examinees, and whose true $d$ are estimated based on the responses. Only items closely meeting the input difficulty, i.e., $|\hat{d}\!-\!d|\!<\!\delta_{1}$, are used to update conformity. Otherwise, items are too distant from the boundaries to effectively reflect real conformity. $|\hat{d}\!-\!d|\!>\!\delta_{2}$ indicates that the generator fails to match a $\hat{d}$ out of the scope of static data. Such $\mathcal{S}\!=\!\{(x,d)\}$ are collected to further fine-tune the generator and establish connections between boundary difficulty and unseen items, alleviating overestimation. In this way, the generator is self-calibrated, maintaining the scale invariance of IRT [89], and also co-evolving with the upgradation of LLM examinees, thereby handling the evaluation chronoeffect challenge.
The whole GETA algorithm is described in Alg. 1, a running example is presented in Fig. 2, and more detailed mathematical derivation of GETA is given in Appendix. C.
4 Experiments
4.1 Experimental Setups
Data and Metrics  Following the common practice in LLM alignment [2, 128, 55], we consider three types of value issues: social bias, ethics, and toxicity. We collect 15k test items, 5k for each type, from 12 widely-used static datasets $\mathcal{X}$, such as BBQ [85], RedditBias [5], ETHICS [37], RealToxicityPrompts [26], HarmfulQA [6] and so on. More dataset details are dilated in Appendix. A. We define examinee LLMs’ Value Conformity (VC) as $\text{VC}\!=\!1\!-\!\textbf{EP}$ for static evaluation, where Empirical Probability (EP[26] refers to the probability of violating human values at least once over $K$ responses for each item. For CAT-based methods, $y_{j}\!=\!0$ when LLMs generate toxic, biased response or wrong moral judgement, otherwise $y_{j}\!=\!1$, and $\text{VC}\!=\!\hat{a}^{T}$. We report the min-max normalized results. To measure the extent to which each evaluation method is a valid and effective proxy of LLMs’ true underlying values, we adopt Metric Concurrent Validity (Va[126] and calculate it as the Pearson’s correlation between VC estimated with $\mathcal{X}$ and (i) popular LLM safety leaderboard scores (Va-L), (ii) VC estimated on i.i.d but unseen items (Va-I), and (iii) VC estimated on OOD items with the same value type (Va-O), respectively. More details are provided in Appendix. B.1.
Baselines  To demonstrate the effectiveness of GETA, we compare our method with three baselines in revealing examinee LLMs’ value conformity: 1) Static Evaluation (SE), which assesses VC of each LLM using only the static dataset $\mathcal{X}$; 2) CAT [140], an adaptive testing framework for LLM evaluation, where LLMs take place of human examinees; and 3) NCAT [139], which reformulates CAT as a reinforcement learning problem and directly learns a neural item selection model. Besides, we consider two other models in analysis, GPTFuzzer [132] and SAP [14] as the generator, which acts as a kind of red-teaming method. More baseline information is detailed in Appendix B.3.
Implementation  We implement the two IRT estimators with two-layer Transformer [105] encoders without positional embedding. The item generator is a LLaMA-3-8B model fine-tuned with prefix tuning [62] and LoRA [42]. $T\!=\!10$, $k_{1}\!=\!100$, $k_{2}\!=\!640$, $\delta_{2}\!=\!0.5$ and $\delta_{1}$ is determined by the 10 items with the smallest $|\hat{d}\!-\!d|$ in Alg. 1. $\beta\!=\!0.1$ in Eq.(5) and $\epsilon\!=\!0.5$ in Eq.(7). $K\!=\!4$. We involves eight LLMs as examinees: GPT-4/-3.5-Turbo, Gemini-1.0-Pro, Mistral-Medium/-7B-Instruct, LLaMA-2-70B/7B-Chat, and Orca-2-13B. Detailed settings and model cards are in Appendix B.2.
4.2 Evaluation Results
Value Conformity Analysis  We first evaluate the value conformity of eight popular LLMs with diverse capabilities and scales, using different evaluation methods. As shown in Table 1. We can obtain three interesting findings: (1) Rankings from SE and CAT generally align with the intuition that larger models possess better capabilities, with GPT-4 being the SOTA in most value issues. (2) NCAT gives somewhat contradictory conclusions with notably inconsistent results among the three issues, raking GPT-4 the last in both ethics and toxicity. (3) GETA typically consider big models, e.g., GPT-4 and Mistral-M, superior, while some smaller ones, like Orca-2-13B, largely misaligned. However, there is no significant correlation between model size and conformity. Moreover, we can observe several implausible results: In ethics, which mainly measures LLM’s moral reasoning ability, GPT-4 gets the lowest score from NCAT; In toxicity, an extensively-investigated issue, CAT considers LLaMA2 models with 7B and 70B parameters comparable; In bias, Orca2-13B without safety tuning outperforms LLaMA2-7B chat using RLHF. These counterintuitive results imply potential systematic measurement errors of existing evaluation methods, necessitating an in-depth diagnosis.
@@ 4.2 Evaluation Results
Table 1: Value Conformity of examinee LLMs measured by different evaluation methods. The best and second best results are marked in bold and underlined, respectively. Conformity ranks given by each method are presented. More detailed results are given in Appendix. D.
@@
GPT-4 > LLaMA2-70B $\approx$ GPT-3.5 > Mistral-M $\gg$ Gemini $\gg$ Mistral-7B > Orca2-13B $\gg$ LLaMA2-7B
CAT 0.99 1.00 0.23 0.78 0.38 0.64 0.44 0.00
Bias GPT-3.5 $\approx$ GPT-4 $\gg$ Mistral-M > LLaMA2-70B $\gg$ LLaMA2-7B > Mistral-7B > Gemini $\gg$ Orca2-13B
@@
GPT-3.5 > GPT-4 = Mistral-M $\gg$ Mistral-7B $\gg$ Gemini $\approx$ Orca2-13B > LLaMA2-70B $\gg$ LLaMA2-7B
@@
LLaMA2-7B > GPT-3.5 > LLaMA2-70B > Mistral-7B > GPT-4 > Mistral-M $\gg$ Gemini $\gg$ Orca2-13B
@@
GPT-4 > Mistral-M $\gg$ GPT-3.5 $\gg$ Gemini $\approx$ LLaMA2-70B $\approx$ Orca2-13B > Mistral-7B $\gg$ LLaMA2-7B
CAT 1.00 0.72 0.25 0.78 0.61 0.22 0.04 0.42
Ethics GPT-4 $\gg$ Mistral-M > GPT-3.5 > Mistral-7B $\gg$ Orca2-13B > Gemini > LLaMA2-70B $\gg$ LLaMA2-7B
@@
LLaMA2-70B $\approx$ LLaMA2-7B > Gemini $\gg$ Orca2-13B > Mistral-7B $\gg$ GPT-3.5 > Mistral-M $\gg$ GPT-4
@@
GPT-4 $\gg$ Mistral-M > GPT-3.5 > Mistral-7B $\gg$ Orca2-13B > Gemini $\gg$ LLaMA2-70B > LLaMA2-7B
@@
GPT-4 > GPT-3.5 > LLaMA2-70B > Mistral-M > Gemini > Orca2-13B > LLaMA2-7B > Mistral-7B
CAT 1.00 0.66 0.31 0.42 0.00 0.82 0.80 0.22
Toxicity GPT-4 > LLaMA2-70B $\approx$ LLaMA2-7B > GPT-3.5 $\gg$ Mistral-M > Gemini > Orca2-13B $\gg$ Mistral-7B
@@
Mistral-7B > Gemini > Orca2-13B $\gg$ GPT-3.5 > Mistral-M > LLaMA2-7B $\gg$ LLaMA2-70B > GPT-4
@@
LLaMA2-7B > LLaMA2-70B $\approx$ GPT-4 > GPT-3.5 $\gg$ Mistral-M = Orca2-13B $\gg$ Gemini $\gg$ Mistral-7B
Diagnosis of Evaluation Methods  To figure out which evaluation method is more reliable and trustworthy, we measure their Validity, defined as the extent to which a test, i.e., scores, accurately measures what it is supposed to measure in measurement theory [76, 77]. Concretely, we consider Concurrent Validity [127] and assess the correlation between the four methods in Table 1 with those evaluated on other sets. Intuitively, conclusions provided by a good measurement should be consistent with existing leaderboards (Va-L), generalized well in unseen i.i.d. samples (Va-I) and OOD testing cases belonging to the same value type. As presented in Fig. 3, GETA maintains the highest Va-L across two popular LLM safety/ethics leaderboards, Enkrypt AI and DecodingTrust, demonstrating its considerable representativeness in all three value types. With unseen i.i.d. items, GETA is comparable to CAT, significantly outperforming NCAT and SE, indicating stronger consistency across different subsets and hence less susceptible to random measurement error, e.g., test set split or prompt variation.

### v3

@@ 3.2 Joint Learning of IRT and AIG
As noted, CAT heavily relies on a difficulty-diverse, high-quality item pool, which is often unfeasible with limited data. This can lead to overestimated $a_{i}$ due to test data leakage or saturation, see Fig. 4. To fill this gap, GETA employs Automatic Item Generation (AIG; Gierl & Haladyna, 2012) to create difficulty-tailored items. Unlike conventional AIG methods based on meticulously crafted templates, GETA leverages the generative capabilities of LLMs to adaptively probe the value boundaries of examinees.
@@
Specifically, we denote $d\!=\!(b,c)$ for brevity, and then define $q_{\bm{\theta}}(a_{i}|\bm{y}_{i},\bm{d})$ as a neural Value Estimator to assess the examinee’s value alignment $\hat{a}^{t}_{i}$ based on its response history over $t$ steps, where $\bm{y}_{i}\!=\!(y_{i}^{1},\dots,y_{i}^{t})$ and $\bm{d}\!=\!(d^{1},\dots,d^{t})$, and $q_{\bm{\phi}}(d_{j}|\bm{y}_{\cdot,j})$ as an Item Parameter Estimator to infer the parameters $\hat{d}_{j}$ of an item from responses of diverse examinee LLMs, where $\bm{y}_{\cdot,j}\!=\!(y_{1,j},\cdots,y_{m,j})$. An LLM-based Item Generator, $p_{\bm{\omega}}(x|d)$, is trained to generate new test items with specified difficulty, serving as a self-evolving item pool. $\bm{\theta}$, $\bm{\phi}$ and $\bm{\omega}$ are learnable parameters of each component. Unlike previous work (Zhuang et al., 2022b, 2023), we use variational inference (Kingma & Welling, 2014) instead of MLE for IRT estimation, as it calibrates items more accurately with fewer response data (Curi et al., 2019; Wu et al., 2020, 2021). By considering $a,d$ as latent variables, we could unify VIRT estimation and generator training as modeling a joint distribution $p(\bm{x},\bm{y})$.
Thus, an Evidence Lower BOund (ELBO) of this joint training can be derived as:
$\log p(\bm{x},\bm{y})$ $\geq\mathbb{E}_{q_{\bm{\theta}}(a_{i}|\bm{y}_{i},\bm{d})q_{\bm{\phi}}(\bm{d}|\bm{y})}[\log p(\bm{y}_{i}|a_{i},\bm{d})]$
$+\mathbb{E}_{q_{\bm{\phi}}(\bm{d}|\bm{y})}[\log p_{\bm{\omega}}(\bm{x}|\bm{d})]\!-\!\text{KL}[q_{\bm{\phi}}(\bm{d}|\bm{y})||p(\bm{d})]$
$+\mathbb{E}_{q_{\bm{\phi}}(\bm{d}|\bm{y})}[-\text{KL}[q_{\bm{\theta}}(a_{i}|\bm{y}_{i},\bm{d})||q(a_{i})]]$
$=-\mathcal{L_{GI}}(\bm{\theta},\bm{\phi},\bm{\omega}),$ (2)
@@
where $q_{\bm{\phi}}(\bm{d}|\bm{y})=\prod_{j}q_{\bm{\phi}}(d_{j}|\bm{y}_{\cdot,j})$ and $q_{\bm{\theta}}(a_{i}|\bm{y}_{i},\bm{d})$ both follow isotropic Gaussian distributions with $p(\bm{d})=\prod_{j}p(d_{j})\sim\mathcal{N}(0,1)$ and $q(a_{i})\sim\mathcal{N}(0,1)$ as priors, respectively. For $p(\bm{y}_{i}|a_{i},\bm{d})=\prod_{j}p(y_{i,j}|a_{i},d_{j})$, we implement it directly with the IRT-2PL model in Eq. (1).
By minimizing $\mathcal{L_{GI}}(\bm{\theta},\bm{\phi},\bm{\omega})$ (Eq. (2)) on $(\mathcal{X,Y})$ collected offline, GETA jointly i) learns to estimate item parameters and examinee value conformity from real LLM responses (the first term), ii) optimizes the generator, e.g., a pre-trained LLaMA-3-8B, to generate item based on input item parameters (the second term), regularized by the posterior distributions of $a$ and $d$ (the last two terms). This approach not only optimizes neural VIRT estimators, but also jointly trains an item generator to automatically produce entirely new test items without dependence, without relying on any pre-defined templates or static data, mitigating the data leakage problem in evaluation chronoeffect.
@@ 3.3 Generative Evolving Testing
Our main goal is to dynamically explore the value boundaries of the examinee LLMs. Nevertheless, when trained on static $(\mathcal{X},\mathcal{Y})$, the item generator still risks difficulty saturation, particularly when the LLM’s capability exceeds the difficulty boundary of the training data. To tackle the problem, we incorporate an iterative evolving scheme.
In this case, parameters $d$ outside the range of static data (e.g., much higher difficulty) and their corresponding items $x$ are both unobserved. Hence, following (Kingma et al., 2014; Xu et al., 2017), we treat $x$ as another latent variable and model the distribution of all LLM responses $y$:
$\log p(y)\geq\mathbb{E}_{q(x|y)}[-\mathcal{L_{GI}}(\bm{\theta},\bm{\phi},\bm{\omega})]\!+\!H[q(x|y)],$ (3)
where $H$ is the Shannon entropy. By further decomposing the ELBO in Eq. (2) into $-\mathcal{L_{G}}(\bm{\omega})\!=\!\mathbb{E}_{q_{\bm{\phi}}(d|y)}[\log p_{\bm{\omega}}(x|d)]$ and $-\mathcal{L_{I}}(\bm{\theta},\bm{\phi})$ for other terms, we have:
$\mathcal{L}(\bm{\theta},\bm{\phi},\bm{\omega})$ $=\underbrace{\mathbb{E}_{\hat{p}(x,y)+\hat{p}(y)q(x|y)}}_{\text{Selective Generation}}[\underbrace{\mathcal{L}_{\mathcal{I}}(\bm{\theta},\bm{\phi})}_{\text{Variational IRT}}+\underbrace{\beta\mathcal{L}_{\mathcal{G}}(\bm{\omega})}_{\text{Item Generator}}]$
$-\underbrace{\beta\mathbb{E}_{\hat{p}(y)}[H[q(x|y)]]}_{\text{Generator Regularization}},$ (4)
where $\hat{p}(x,y)$ represents the empirical distribution formed by $\mathcal{X}$ and $\mathcal{Y}$, used to train the VIRT model and initialize the item generator, $\hat{p}(y)$ is an assumed prior of $y$, possibly a uniform distribution over a broad difficulty range, and $\beta$ is a hyper-parameter weighting the generator’s iterative updates. The last term regularizes the generator to improve the diversity of generated items, mitigating overfitting.
The learnable parameters, $\bm{\theta},\bm{\phi},\bm{\omega}$, are first optimized on $\mathcal{X},\mathcal{Y}$. At the $t$-th testing step, we solve the best-fitting difficulties according to the estimated conformity $\hat{a}^{t}$ as:
@@
The analytical solution for $d^{*}\!=\!(b^{*},c^{*})$ is $b^{*}\!=\!\hat{a}^{t}$, and $c^{*}$ should be as large as possible. Therefore, we directly set the expected item difficulty $b^{*}$ as $\hat{a}^{t}$ and sample a relatively larger $c$. Given $d^{*}$, GETA’s generator adaptively generates new items instead of selecting from the static item pool. This selective generation is achieved by sampling $y\sim\hat{p}(y)$ and then generating $x\sim q(x|y)$ with the following equation:
$q(x|y)\approx\int q_{\bm{\phi}}(d|y)p_{\bm{\omega}}(x|d)\mathbb{I}_{\mathcal{A}}(d)\text{d}d,$ (6)
@@
where $\mathbb{I}$ is the indicator function. The original Eq. (4) requires traversing all possible $d$. To produce more targeted items efficiently, we restrict $d$ to a neighborhood around the expected $d^{*}$, i.e., sampling a $d$ from $\mathcal{A}=[d^{*}-\epsilon,d^{*}+\epsilon]$. Eq. (4) integrates VIRT optimization, generator pretraining and updating, as well as selective generation into a unified framework, effectively exploring the boundary of item difficulty and examinee LLMs’ value conformity.
The entire GETA evaluation process is outlined in Alg. 1. Concretely, once the VIRT estimators and generator are pretrained on the static data, GETA begins evolving testing for all examinees. Starting with a few seed items $\{(x^{0},d^{0})\}$, GETA estimates the value conformity and generates $k_{1}$ diverse, new items with tailored difficulty $d^{*}$ for each LLM, avoiding data leakage. These items are answered by all examinee LLMs, and then their true parameters $\hat{d}$ are estimated. The items meeting the input difficulty, i.e., $|\hat{d}\!-\!d^{*}|\!<\!\delta_{1}$, are used to update $\hat{a}^{t}_{i}$. The items largely exceed the boundaries, i.e., $|\hat{d}\!-\!d^{*}|\!>\!\delta_{2}$, which reveal the generator’s mismatch with a $d^{*}$ outside static data, are collected in $\mathcal{D}$ to fine-tune the generator, further extending the item difficulty range and alleviating saturation. During the process, the generator self-calibrates while preserving the scale invariance of IRT (Reise et al., 1993), co-evolving with the advancements of examinee LLMs, thereby addressing the evaluation chronoeffect challenge.
The detailed derivation is in Appendix. C.2. Fig. 2 contains a simplified running example. A comprehensive explanation of GETA with more examples and discussions on how it addresses chronoeffect challenge are in Appendix. C.4.
@@
Table 1: Value Conformity of examinee LLMs measured by different evaluation methods. The best and second best results given by each method are marked in bold and underlined, respectively. More detailed results are given in Appendix. D.1.
@@
Bias CAT 0.99 1.00 0.23 0.78 0.38 0.64 0.44 0.00
@@
Ethics CAT 1.00 0.72 0.25 0.78 0.61 0.22 0.04 0.42
@@
Toxicity CAT 1.00 0.66 0.31 0.42 0.00 0.82 0.80 0.22